The seismic activity has continued to increase during the last 24 hours but is still at a low level. About forty long-period earthquakes were recorded by most of the stations in the network, along with several signals generated by rockfalls on the dome.
The northern and eastern EDM triangles were measured today. The results show only small changes in the line lengths since the triangles were last measured.
Visual observations were made from the helicopter this morning. No major changes since yesterday were observed. The top of the dome was at a height of 2860 feet yesterday, and had grown vertically by about 100 ft over the last 10 days. The dome is still confined within the crater.
Gas and particle samples were collected at the Gages seismic station, at Upper Amersham and from the helicopter. The samples collected will enable full characterisation of the volcanic plume, which continues to be emitted at a high rate and includes steam as well as volcanic gases and particles.